CH/Doc/93/11

Visa Details

Experimental Work in Australia

Author: R.P. Ward

Date: 6 July 1993

Procedure for Obtaining an Australian Visa

1. Everyone (staff and students) should apply for a Visiting Academic Visa (class 419) on form 147.

2. The completed form should be submitted to an office of the Australian Government, along with a cheque for £55 payable to the “Commonwealth of Australia”, three passport photographs (signed on the reverse) and your letter of invitation to the ANU.

3. I recommend that applications be made in person rather than by post. The issue of this visa takes around ten minutes and is very straightforward. It is quite acceptable to submit an application on behalf of someone else. Thus, I suggest that one person from each institution perform the task on behalf of their colleagues.

4. If are resident in Southern England, apply to:

Australian High Commission Australia House Strand London WC2B 4LA Tel: 071–379 4334

If you are resident in the Midlands or Northern England, apply to:

Australian Consulate Chatsworth House Lever Street Manchester M1 2DL Tel: 061–228 1344 5. My experience of Australian Government Offices is that they are open from

10am till 4pm on weekdays.

6. Holders of Visitor Visas are advised that these are not the correct visas for this sort of trip. Visitor Visas entitle the holder to engage in discussions and attend meetings, but not to perform actual work. You may be able to gain entry to Australia with such a visa, but you run the risk of being refused entry.

7. Visiting Academic visas are issued for a maximum of four months, starting the first day of the month in which the application is made. Thus, if you apply in July your visa will terminate on 1st November; apply in August and it will terminate on 1st December. You must leave Australia before your visa expires.

Application is fairly routine, and need not be carried out too far in advance of the journey. Please get in touch with me if you have any queries.

Roger Ward